The length of a rectangle is 6 inches more than 4 times the width. If the area of the rectangle is 54 square inches, find the length and the width.
Answer by lwsshak3(11628)

The length of a rectangle is 6 inches more than 4 times the width. If the area of the rectangle is 54 square inches, find the length and the width.
***
let x=width of rectangle
4x+6=length of rectangle
width*length=area
x(4x+6)=54
4x^2+6x=54
4x^2+6x-54=0
2x^2+3x-27=0
(2x+9)(x-3)=0
x=3
4x+6=18
width of rectangle=3 inches
length of rectangle=18 inches